Which of the following equations represents an elliptic paraboloid ?

 

 

  1. "z=y^2"
  2. "z=x^2-y^2+2y"
  3. "z=x^2+y^2+2x"
  4. "z=x^2-y^2+2x"
1
Expert's answer
2021-10-11T14:25:54-0400

The equation of an elliptic paraboloid:

"z=Ax^2+By^2"

where A and B have the same sign.


Answer: "z=x^2+y^2+2x"

or "z+1=(x+1)^2+y^2"